# A Survey of Zero-Knowledge Proof Based Verifiable Machine Learning

### v1

@@ III-B3 ZKP based Verifiable Inference
III-C Improvements in ZKML Implementations
Most machine learning algorithms are inherently complex and computationally expensive, making their implementation with existing ZKP solutions highly challenging. This complexity often renders direct proof generation and verification infeasible for many practical applications. Specifically, there are two primary challenges in implementing ZKML:
  • Generality Limitations: Most machine learning tasks rely on floating-point computations and nonlinear operations. However, ZKP, as a cryptographic technique, operates on linear arithmetic over finite fields. Floating-point and nonlinear operations in ML cannot be directly translated into the arithmetic circuit representations required by ZKP over finite fields. As a result, ZKP algorithms must be specifically tailored to handle the floating-point computations and nonlinearities of different ML models. For example, the ReLU activation function, $f(x)=max(x,0)$ used in neural networks, though computationally simple in conventional ML implementations requires specialized adaptations in ZKML implementations. ReLU inherently requires inequality comparisons ($x>0$) and conditional branching, which cannot be natively executed in finite field arithmetic. To circumvent this, ReLU can be expressed as $x\cdot\mathbb{I}(x{\geq}0)$ by introducing auxiliary binary variables for the indicator function $\mathbb{I}$ (as proposed in [32]). Alternatively, ReLU can be approximated using polynomial functions. Moreover, these ReLU-specific adaptations fail to generalize to other activation functions. This demonstrates how ZKML systems sacrifice generality for functionality—each unique nonlinear operation demands custom constraint system design, precision-robustness tradeoff analysis, and specialized optimization passes.
  • Efficiency Barriers: Modern machine learning models, such as deep neural networks, typically involve a large number of parameters, resulting in massive arithmetic circuit representations. Consequently, the proof generation and verification processes in ZKP can become prohibitively time-consuming. For instance, Groth16, one of the most widely used zk-SNARK protocols, has a proof generation complexity of $O(|C|E)$, where $|C|$ denotes the size of the arithmetic circuit (directly proportional to the computational workload), and $E$ represents the cost of group exponentiation. To prove the inference using a deep neutral network model like VGG16 that has 138 million parameters, the value of $|C|$ and $E$ can be calculated as follows. Each convolutional layer requires round 5,000 multiplication gates per parameter due to weight-input multiplications and activation function approximations; total circuit size reaches $|C|\approx 7.2\times 10^{11}$ gates (720 billion gates); Groth16 proof generation requires 2 group exponentiations per multiplication gate, leading to $E=1.44\times 10^{12}$ exponentiations. For a VGG16-sized circuit using BN254 elliptic curve (254-bit base field): the evaluation key size is $|C|\times 3\times 64~{}/10^{12}=(720\times 10^{9}\times 192)/10^{12}=138.24$ TB; the verification key size is $(|C|+2)3.64~{}/10^{12}=(720\times 10^{9}\times 192)/10^{12}=46.08$ TB; the proof generation time is over 30 years on a 256-core AWS c6i.32xlarge instance. This demands an extraordinary amount of memory and computational resources. Thus, efficiently generating proofs for ML computations remains a significant challenge.
In the following, we explore how existing research efforts tackle the challenges of improving the generality and efficiency of ZKML implementations. These efforts are categorized into three main areas: enhancing generalization, improving efficiency, and addressing other aspects across various types of ZKML, as summarized in Fig. 7.
III-C1 Improvement of Generality
To bridge the gap between floating-point machine learning operations and finite-field ZKP primitives, the key lies in converting decimals into integers through quantization functions that minimize precision loss while maintaining model performance. This process also involves determining the optimal number of variable bits. On one hand, the range of integers representable by a fixed number of bits is inherently limited, potentially causing overflow issues when decimals mapped by the quantization function exceed this range. On the other hand, the computation involving these converted integers must retain as many significant bits as possible to reduce the impact of diminished computational precision on overall model accuracy.
To ensure accuracy, in SafetyNets [34], all decimal weights and input data must be within $[-(p-1)/2,(p-1)/2]$. The floating-point parameters $W_{i}^{\prime}$ and $b_{i}^{\prime}$ obtained from training are multiplied by a constant $\beta>1$ and rounded to convert them into integers, i.e., $W_{i}=\lceil\beta W_{i}^{\prime}\rceil$, with the same method applied to inputs using a scaling factor $\alpha$, i.e., $x=\lceil\alpha x^{\prime}\rceil$. To ensure isotropic scaling of all values in the network, set $b_{i}=\lceil\alpha z^{l-1}\beta(z^{l-1}+1)b_{i}^{\prime}\rceil$. This method achieves an integer computation process. In VeriML [29], each input is restricted to at most $l$ decimal places by multiplying the input by $2^{l}$ to convert it to an integer, applying appropriate scaling factors in the equations. zkCNN [10] adopts the quantization technique proposed by Jacob et al. [72] to encode floating-point numbers as integers. This quantization scheme is an affine mapping from integer $q$ to real number $a$. Specifically, $a=L(q-Z)$, where the quantization parameter $L$ is a real number called the quantization scale, and $Z$ is an integer called the quantization zero point. Using quantization, each value of data samples and model parameters is represented as a Q-bit integer $q$, allowing addition and multiplication of two real numbers with the same scale to naturally represent integer addition and multiplication. Kang [39] employs quantized DNNs to avoid the costly overhead of floating-point arithmetic, representing weights and other floating-point parameters as 8-bit integers, similar to the method in zkCNN. This quantization scheme is an affine mapping from integer $W_{\text{quant}}$ to real number $W,w=(W_{\text{quant}}-z)\cdot(a/b)$, where $W_{\text{quant}}$ and $z$ are 8-bit integer weights and zero point, and $a/b$ represents fixed-point approximation. Choosing lower precision values for $a$ and $b$ leads to slight accuracy degradation, but significantly improves proof and verification performance. However, the aforementioned methods are only applicable to problems within specific ZKP systems and lack generality. The above methods are only applicable to problems within specific zero-knowledge proof systems. Next, existing research solutions that effectively address generality issues will be introduced, which can be used directly or indirectly in arithmetic processing in the future.
Figure 7: The categorization of ZKML works based on improvements in generality, efficiency, and other aspects.
To further address the issue of generality, enabling future direct or indirect use in arithmetic processing, Weng et al. proposed the Mystique system [13], which allows efficient conversion between arithmetic and Boolean values, between public commitments and private authenticated values, and between fixed-point and floating-point numbers. The authors also design an efficient zero-knowledge proof for matrix multiplication, making the number of private multiplications required sublinear in matrix size. Notably, this system has been integrated into the privacy-preserving framework Rosetta[73] based on TensorFlow[74], meaning developers can directly call this system and ignore the cryptographic details involved. ZEN [38] proposed a new quantization algorithm that avoids floating-point zero-knowledge proofs, including two R1CS-friendly optimizations: sign bit grouping and remainder-based verification. Compared to existing quantization works, ZEN saves 73.9 times R1CS on convolution kernels and 8.4 times R1CS on fully connected kernels without causing additional accuracy loss. Additionally, ZEN can convert a floating-point PyTorch model into an arithmetic circuit over a finite field. Kang et al. [39] used the zero-knowledge proof scheme Halo2 to customize an ImageNet-scale zk-SNARK circuit for the MobileNet v2 model. To address the high verification cost of division operations in the circuit, two custom gates were designed for linear layers to represent division. For nonlinear layers, lookup parameters were used to reduce the representation cost of division. Three protocols based on this scheme were also proposed for verifying model accuracy, prediction results, and search term-predicate matching. Compared to other works (including ZEN [38], vCNN [37], pvCNN [41], zkCNN [10]), this scheme improves proof time on MobileNet by at least ten times.
The above works focus on addressing the challenges of quantization techniques for floating-point operations in zero-knowledge proof systems, providing solutions to bridge the gap between machine learning computations and finite-field arithmetic. However, an additional challenge arises when adapting widely used machine learning models, particularly neural networks, to these systems. While operations such as matrix multiplication and convolution in neural networks can be effectively represented using basic addition and multiplication, most activation functions are inherently nonlinear and cannot be directly expressed within arithmetic circuits.
Currently, methods for handling activation functions in neural networks include: SafetyNets [34], VeriML [29], zkMLaaS [31], zkCNN [10], and zkDL [32]. For SafetyNets [34], it uses a quadratic activation function. This directly avoids complex computations but may impact network performance due to poor activation functions. For VeriML [29], it uses Remez approximation to approximate the sigmoid function and quadratic functions to approximate the ReLU function, thereby addressing the issue caused by the poor activation in SafetyNets. For zkMLaaS, it utilizes the method proposed by Sav et al. [73] to minimize the least squares error when approximating activation functions. However, this method also leads to performance degradation due to approximation. For zkCNN [10], it uses a bit decomposition method to directly compute the ReLU function in segments. Fan constructs an operation matrix where the input of the ReLU activation layer is Hadamard-multiplied with the operation matrix to obtain the output of the ReLU layer. This method has minimal impact on model performance but introduces additional computation and design overhead. For zkDL [32], it further proposes a zero-knowledge proof protocol optimized for ReLU activation functions, the zkReLU protocol, which introduces auxiliary inputs to verify the forward and backward propagation processes involving ReLU, retains the tensor structure of deep learning, and combines zero-knowledge inner product proofs with the GKR protocol designed for optimized arithmetic circuits, achieving significant improvements in proof generation speed and size reduction.
III-C2 Improvement of Efficiency
Recent advancements in ZKP systems for machine learning have focused on improving efficiency and scalability through innovative algorithmic designs and circuit optimizations. These techniques aim to bridge the gap between the computational complexity of machine learning models and the constraints of ZKP frameworks by leveraging methods such as matrix representations, efficient arithmetic circuit constructions, and tailored proof strategies. Key ideas include transforming complex operations like convolutions and nonlinear activation functions into simpler arithmetic forms, optimizing proof generation with probabilistic algorithms, and modularizing model components to streamline verification. Additionally, techniques like random sampling, table lookups, and specialized commitment schemes have been introduced to reduce proof costs and enhance scalability. These foundational ideas set the stage for the following research efforts, which tackle specific challenges in validating neural networks, decision trees, and other machine learning models within ZKP systems.
In SafetyNets [34], the GKR protocol is designed based on matrix multiplication equations to achieve efficient verification of neural network computations. Unfortunately, for activation functions and pooling layers, SafetyNets [34] can only support specific quadratic activation functions and sum pooling, making it relatively impractical. To address this issue, Liu et al. proposed zkCNN [10], a zero-knowledge proof scheme for convolutional neural networks based on the GKR protocol, which improves the efficiency of zero-knowledge proofs by optimizing the proof cost of convolution computations in convolutional neural network models. Specifically, zkCNN [10] designs a new GKR protocol for verifying fast Fourier transform (FFT) computations and constructs a protocol for two-dimensional convolutions based on FFT, achieving an additional proof time complexity of $O(n^{2})$, faster than the convolution computation itself. In terms of performance, zkCNN is 11.2 times and 213 times faster than vCNN and ZEN on LeNet, respectively. Meanwhile, Fan [42] et al. also focused on convolution computation, converting the computation into simple arithmetic expressions in matrix form. For convolution layers, the im2col method is used to represent 3D convolutions as 2D matrices, thus converting convolution computation into equivalent matrix multiplication. Pooling layers also use the im2col method to reduce 3D data to 2D representations. Activation functions ReLU and Softmax are also expressed in the form of matrix multiplication. ReLU is represented as the input matrix multiplied by a matrix with elements 0 or 1, while Softmax is represented as the output matrix multiplied by a vector of exponentials. Additionally, all matrix computations in convolutional neural networks can be optimized using Freivalds’ algorithm [75], significantly improving the efficiency of setup and proof generation.
Zhan [52] leverages depthwise separable convolutions, Hesamifard’s approximate activation function scheme [76], and adaptive average pooling based on specified output sizes to transform convolutional layers, normalization layers, activation layers, and pooling layers into simplified linear arithmetic expressions, while optimizing them according to circuit-specific characteristics. The overall network structure of the large language model is modularized into three components: the head, the backbone, and the tail. To manage the connections between these components, a hash function is introduced, ensuring efficient and seamless integration.
Theoretically, compared to works like SafetyNets [34] and VeriML [29], vCNN [37] makes certain improvements in the proof of convolution computation. vCNN [37] extends the original zk-SNARKs scheme based on quadratic arithmetic programs (QAPs) to zk-SNARKs based on quadratic polynomial programs (QPPs) by leveraging the properties of convolution computation. vCNN [37] uses polynomial expressions of vectors to make the proof of convolution computation more adaptable and efficient. For pooling and activation layers, vCNN [37] retains the QAP-based zk-SNARK and generates continuity proofs connecting adjacent layers through Committed Proof SNARK (CP-SNARK). Thus, QAP and QPP-based zk-SNARKs prove the correctness of intra-layer computation, while CP-SNARK proves the continuity of each layer’s computation, ultimately proving the correctness of the entire convolutional neural network computation.
Inspired by vCNN [37], Weng et al. further proposed pvCNN [41], a framework for verifiable convolutional neural networks. They proposed a zk-SNARKs scheme based on quadratic matrix programs (QMP) on top of the QPP method in vCNN. pvCNN [41] extends the representation capability of wires in the circuit from arrays to matrices and reduces the number of multiplication gates in convolution operations, thus reducing the circuit size and improving efficiency. Additionally, because neural networks are layered, multiple proofs for different inputs of the same CNN layer can be merged into one proof. In terms of performance, this scheme is theoretically compared with SafetyNets[34], zkCNN [10], and vCNN [37], showing significant improvement in proof time over the aforementioned works. Experimental results also indicate that QMP-based zk-SNARKs are more efficient in convolution operations than QAP-based methods.
zkLLM [46] introduces tlookup, which allows parallel lookup of parameters in non-arithmetic tensor operations in deep learning, providing a solution without asymptotic overhead. Lu [47] discards bit decomposition in generating proofs for nonlinear layers, instead converting nonlinear relationships in neural networks into range and exponential relationships, completing the proof of nonlinear layers through range proofs and lookup proofs, achieving about two orders of magnitude reduction. Hao [50] uses table lookup for ZK proofs of nonlinear functions, addressing the challenge of large tables by decomposing inputs with large bit lengths into a fixed number of smaller numbers to obtain smaller tables, and designs modules for number decomposition, comparison, and truncation to effectively utilize table lookup. Compared to Mystique [13], the protocol achieves a 50 to 179 times improvement in runtime for widely used nonlinear functions in machine learning, such as ReLU, sigmoid, and GELU.
For decision tree models, to verify the output, the prover must first commit to the decision tree and then prove the validity of the verification path to the verifier. However, converting each comparison on the verification path into an arithmetic circuit is very costly. To address this issue, ZKDT [32] reduces the cost of proof generation by inserting designed sibling nodes on the verification path. Singh et al. proposed a zero-knowledge verifiable scheme for distributed AI pipelines, including a privacy-preserving verification scheme for decision tree inference [40]. The distributed AI pipeline assigns different steps of data collection, model training, and model prediction to independent participants, such as data owners, model owners, and model users. Compared to ZKDT [32], this scheme avoids costly hash operations by changing the representation and commitment of decision trees. Moreover, by improving access methods in arithmetic circuits, it reduces the access cost of different operations in prediction path verification, further reducing the number of multiplication gates in arithmetic circuits. For decision tree inference tasks, the circuit complexity generated by this scheme is ten times more efficient than zkDT [32]. Lookup arguments [44] encode decision trees as a matrix, and by committing to this matrix encoding, the decision tree is committed. Due to the efficiency characteristics of matrix lookup parameters, the prover’s time complexity is independent of the decision tree size, improving proof time by nearly an order of magnitude and verification time by two orders compared to ZKDT.
The goal of improving proof efficiency by optimizing algorithm structure is to reasonably partition the algorithm and apply appropriate proof strategies to proportionally enhance proof efficiency. For example, verifying only a portion of the rounds during the training process can reduce the number of proofs generated. Since the same process is iterated multiple times during training, VeriML chooses to use a few rounds of the training process as a challenge for verification, thereby reducing proof costs. By pre-storing inputs and outputs of some iterations during training and committing to them, the prover can retrieve specified iterations and generate proofs of their computation process as required by the verifier. Some small optimizations are proposed for the six machine learning models supported by VeriML to improve proof efficiency. Additionally, VeriML uses an on-chain protocol to protect the confidentiality of trained models for fair transactions.
Hydra [30] uses the GKR protocol to compute sub-circuits split from the original circuit. Through a polynomial commitment scheme, Hydra simplifies the proof process. Meanwhile, since sub-circuits are independent of each other, they can be generated and verified in parallel, improving efficiency. Additionally, a bottom-up quantization algorithm is proposed to reduce the impact of integration on accuracy. Compared to SafetyNets, the Hydra protocol improves efficiency on neural networks by four times.
Lycklama [51] also addresses the challenges in commitment schemes, proposing two novel methods, Apollo and Artemis, to mitigate the significant overhead associated with committing to models and data. Apollo streamlines the construction process by minimally adapting the arithmetization of the witness within the SNARK framework, whereas Artemis is designed to support any homomorphic polynomial commitment scheme. For the VGG model, this work achieves a substantial reduction in the overhead of commitment checks, lowering it from 11.5x to 1.2x.
zkMLaaS [31] focuses on addressing the issue of input data volume and adopts the idea of random sampling. By randomly selecting and challenging committed epochs and iteration numbers, proof costs are proportionally reduced, similar to VeriML. Regarding convolution operations in CNNs, the optimization idea is similar to the method proposed by Fan et al. [42]. The im2col algorithm is applied to convert convolutions into matrix multiplications, and Freivalds’ algorithm is further used to reduce the overhead of matrix multiplication. Compared to directly using zk-SNARKs, zkMLaaS saves approximately 273 times the proof overhead.

### v2

@@ III-B3 ZKP-based Verifiable Inference
III-C Implementation Improvements in ZKML
Despite the rapid progress reviewed above, practical ZKML systems remain difficult to engineer because mainstream ML computation is not naturally aligned with proof systems over finite fields. Two bottlenecks recur throughout the literature:
  • Generality limitations: Most ML models rely on floating-point arithmetic, nonlinear activations, and comparison-style logic, whereas many practical ZKP systems operate over finite-field arithmetic circuits. As a result, common ML operators cannot be translated directly into proof-friendly representations. ReLU, for example, involves comparison and conditional behavior, so it must be encoded through auxiliary binary variables, polynomial approximation, or other operator-specific gadgets. These adaptations rarely transfer cleanly across models or proof systems. In practice, each new activation, normalization rule, or quantization strategy often requires custom arithmetization, precision analysis, and optimization.
  • Efficiency barriers: Even after successful arithmetization, modern ML models induce enormous circuits. Large neural networks lead to high proving time, large commitment and key material, and substantial memory pressure. A naive zk-SNARK treatment of VGG16-scale inference, for instance, would yield an impractically large circuit and correspondingly prohibitive prover cost. Similar bottlenecks arise in witness generation, commitment verification, and proof composition. Reducing end-to-end proof cost without weakening soundness therefore remains a central implementation challenge.
Existing work addresses these issues through two main directions—improving generality and improving efficiency—together with several related system-level optimizations, as summarized in Fig. 7.
III-C1 Improving Generality
A central issue in ZKML is how to bridge real-valued ML computation and finite-field proof arithmetic. The most common first step is quantization: real-valued inputs, activations, and parameters are mapped to integers so that they can be processed inside arithmetic circuits. The design tension is clear. Too few bits increase overflow risk and quantization error, whereas too many bits inflate circuit size and proving cost.
Early systems mainly rely on task-specific quantization strategies. SafetyNets rescales weights and inputs with global constants so that all values remain within the valid field range [46]. VeriML converts fixed-precision decimals into integers by scaling inputs with powers of two [26]. zkCNN adopts affine quantization of the form a=L(qZ)a=L(q-Z), where $L$ is the quantization scale and $Z$ is the zero point [35, 73]. Likewise, the Halo2-based MobileNet v2 implementation in [47] uses low-bit quantized parameters to avoid expensive floating-point arithmetic. These designs make ML computation more compatible with finite-field proof systems, but they remain tightly coupled to particular proof frameworks and model families. Thus, they improve practicality without fully resolving the generality problem.
Figure 7: Taxonomy of representative ZKML studies organized by the main type of improvement they target. The figure first groups prior work into improvements in generalizability, improvements in efficiency, and other system-level directions. Within each branch, representative studies are further organized by verification objective, including verifiable training, verifiable testing, and verifiable inference; inference-oriented work is additionally separated by privacy setting, such as public-input/private-output and private-input/private-output scenarios. This taxonomy highlights how existing ZKML systems differ not only in optimization goals, but also in the ML task and privacy model they support.
More general solutions aim to provide reusable conversion and arithmetization primitives. Mystique offers efficient conversions between arithmetic and Boolean values, between public commitments and private authenticated values, and between fixed-point and floating-point representations, together with an optimized matrix-multiplication proof whose private multiplication cost is sublinear in the matrix size [39]. ZEN introduces quantization-oriented optimizations such as sign-bit grouping and remainder-based verification, allowing floating-point PyTorch models to be compiled more efficiently into finite-field circuits [34]. MSCProof [50] further contributes a modular composition framework based on verifiable evaluation schemes, showing how reusable sumcheck-style components can be composed across sequential ML and image-processing pipelines. The Halo2-based system of [47] similarly uses custom gates and lookup arguments to support ImageNet-scale MobileNet v2. Collectively, these works move ZKML away from purely ad hoc operator rewriting and toward more reusable proof-friendly ML representations and more composable verification pipelines.
Nonlinear operators remain another major source of generality loss. SafetyNets circumvents the issue by restricting networks to quadratic activations and sum pooling [46]. VeriML and zkMLaaS instead rely on polynomial approximation, including Remez-style approximations and least-squares fitting for common activations [26, 27, 74]. zkCNN computes ReLU more directly through bit decomposition, while Fan et al. express ReLU and Softmax in matrix-oriented forms that are easier to verify [35, 49]. zkDL goes further with zkReLU, a protocol specialized for ReLU-related forward and backward propagation that preserves tensor structure while reducing proof cost [28]. More recently, zkGPT avoids full arithmetic emulation of certain nonlinear operators by treating their outputs as advice and verifying them through inverse relations together with range and lookup checks [55]. This progression reflects a broader trend: recent systems increasingly combine approximation, lookup, and advice-based verification in order to balance generality, accuracy, and proof efficiency.
III-C2 Improving Efficiency
Practical ZKML depends not only on expressive arithmetization, but also on keeping prover time, memory footprint, and commitment overhead within manageable ranges. Existing work improves efficiency through several recurring strategies: rewriting ML operators into proof-friendly algebraic forms, exploiting model structure for parallelism or probabilistic checking, and reducing the overhead of commitments and proof composition.
At the operator level, SafetyNets pioneered GKR-based verification for layered neural computations, but its restriction to quadratic activations and sum pooling limited practicality [46]. zkCNN advances this line by designing FFT-aware GKR protocols for two-dimensional convolutions, achieving prover complexity $O(n^{2})$ and substantial empirical speedups over vCNN and ZEN on LeNet [35]. Fan et al. similarly recast convolution, pooling, ReLU, and Softmax into matrix-style computations and then apply Freivalds’ algorithm to reduce setup and proof costs [49]. Related work in [54] further lowers circuit complexity through depthwise separable convolutions, approximate activations, adaptive average pooling, and a modular head–backbone–tail design linked by hash-based integrity checks.
Other systems reduce cost by changing the proof representation of convolutional networks. vCNN extends QAP-based zk-SNARKs to quadratic polynomial programs (QPPs), making convolution proofs more natural while retaining committed-proof links across layers [38]. pvCNN pushes this direction further with quadratic matrix programs (QMPs), which reduce multiplication gates for convolution and enable proof aggregation across repeated CNN layers [33]. For large language models, zkLLM introduces tlookup for parallel verification of non-arithmetic tensor operations, and VeriLoRA transfers this lookup-based philosophy to verifiable LoRA fine-tuning [42, 31].
Efficiency challenges also arise in models beyond CNNs. For decision trees, zkDT reduces path-verification cost through carefully designed sibling information, while later work by Singh et al. and Campanelli et al. further reduces the cost of tree commitment and leaf verification through improved tree encodings and matrix lookup arguments [32, 40, 52]. These advances make proving time less sensitive to tree size and significantly improve verification performance.
At the system level, several works cut overhead by proving less or by composing proofs more effectively. VeriML lowers cost by committing to intermediate states and proving only verifier-selected training rounds [26]. Hydra partitions neural-network training circuits into subcircuits that can be proved and verified in parallel, while also using progressive quantization to limit accuracy loss [45]. Apollo and Artemis reduce the commitment overhead of commit-and-prove SNARK pipelines, bringing the commitment-check overhead for models such as VGG much closer to the cost of the underlying proof system [36]. Finally, zkMLaaS combines random sampling with im2col conversion and Freivalds-based matrix verification, yielding an approximately $273\times$ reduction in proof-generation overhead relative to a more direct zk-SNARK approach [27].